OpenStudy (anonymous):

Tre rolls two standard six-sided number cubes. How many ways can he roll a sum of 7?

OpenStudy (anonymous):

\[nCr = \frac{n!}{r! (n - r)!}\] This is a combination, but in this case it's simpler just to think about it, as there's not many: 1+6=7 6+1=7 2+5=7 5+2=7 3+4=7 4+3=7

OpenStudy (anonymous):

Make sense?